OpenStudy (anonymous):

Is this sum convergent? If so, find the sum. sum_(n=1)^infinity (2^n+n^2+n)/(2^(n+1)n(n+1))

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} {\frac{2^n+n^2+n}{2^{n+1}n(n+1)}} \] is that right?

OpenStudy (anonymous):

Absolutely

OpenStudy (anonymous):

You can divide it into sum of 2 sums : sum 1 : \[\sum_{n=1}^{\infty} {\frac{1}{2n(n+1)}}\] sum 2 : \[\sum_{n=1}^{\infty} {\frac{1}{2^{n+1}}}\]

OpenStudy (anonymous):

converging since the limit of last term ,when n tends to infinity is infinity..

OpenStudy (anonymous):

Thanks anhkhoavo1210, I see how you got that. What would be the next step?

OpenStudy (anonymous):

You see the sum 1 can be divided into : \[\frac{1}{2}(\sum_{n=1}^{\infty} {\frac{1}{n}}-\sum_{n=1}^{\infty} {\frac{1}{n+1}}) \] So you can conclude these sums converge or not?

OpenStudy (anonymous):

Well \[\sum_{n+1}^{\infty} {1 \over n}\] is a harmonic series, which is divergent, correct? I am confused on operating with sums/differences of divergent series, though.

OpenStudy (anonymous):

typo: n+1 should be n=1 in the series notation.

OpenStudy (anonymous):

you know \[\sum_{n=1}^{\infty} {\frac{1}{n+1}}\] can be replaced into \[\sum_{m=2}^{\infty} {\frac{1}{m}}\], is the form of harmonic series?

OpenStudy (anonymous):

I do not really understand that concept. Can you please elaborate?

OpenStudy (anonymous):

put \[m=n+1\]

OpenStudy (anonymous):

Ah, I see. That makes sense.

OpenStudy (anonymous):

But you know : \[\sum_{n=1}^{\infty} {\frac{1}{n}}=1+\frac{1}{2}+\frac{1}{3}+\frac{1}{4}+...\] \[\sum_{m=2}^{\infty} {\frac{1}{m}}=\frac{1}{2}+\frac{1}{3}+\frac{1}{4}+...\] If you use the subtraction here, you have sum 1 = 1/2 That's so surprise. :)

OpenStudy (anonymous):

Ok, I see.

OpenStudy (anonymous):

Correct me if I am wrong: the second series is a geometric series with R = 1/2. Hence the sum of the second series is also 1/2. The sum of the series 1 and series 2 are therefore 1 and the entire series is convergent?

OpenStudy (anonymous):

Of course, the entire series is convergent, but I try to calculate sum 2 x) That is a problem of technic... \[\sum_{k=0}^{\infty} {x^k}=1+x+x^2+...=\frac{1}{1-x}, |x|<1\] I try to find similarities of the above form and sum 2. sum 2 is equivalent to \[\sum_{m=2}^{\infty} {\frac{1}{2^m}}\] , using m=n+1 again Take x=1/2 into the form. The form = \[\sum_{k=0}^{\infty} {(1/2)^k}=1+(1/2)+(1/2)^2+...=1+(1/2)+\sum_{m=2}^{\infty} {\frac{1}{2^m}}\] Then \[\frac{1}{1-(1/2)}=1+(1/2)+\sum_{m=2}^{\infty} {\frac{1}{2^m}}\] Yup it equals to 1/2.

OpenStudy (anonymous):

Thank you! Another question: is there a way to tell from the beginning that the series is convergent?

OpenStudy (anonymous):

sr I don't understand sth what you say. The series here is general?

OpenStudy (anonymous):

How can you tell that the series converges before determining sum?

OpenStudy (anonymous):

of course, use limit and other tests like root test and all..

OpenStudy (anonymous):

In this series above, because you are not sure the convergence of sum 1, you must try to calculate it.

OpenStudy (anonymous):

Awesome. Thanks a lot!